OpenStudy (anonymous):

Suppose that f(x)=3x−6x+5. Use interval notation to indicate where f(x) is concave up or down. Concave up: Concave down:

OpenStudy (hunus):

Is that supposed to be\[f(x)=3x^2-6x+5\]

OpenStudy (anonymous):

Oops! It's suppose to be 3x-6/x+5. Sorry and thanks!

OpenStudy (hunus):

\[f(x)=3x−6/x+5\] \[f'(x)=3+6/x^2\] \[f''(x)=-12/x^3\] \[-12/x^3>0\] \[-1/x^3>0 \] when x<0 \[-12/x^3<0\] when x>0 \[f(x)=3x−6/x+5\] is concave up on the interval \[(-\infty, 0)\] and is concave down on the interval \[(0, \infty)\]

OpenStudy (anonymous):

My online homework said this answer was incorrect :/

OpenStudy (hunus):

:( Are you sure all the information you gave me was correct?

OpenStudy (anonymous):

Uhm Yes, I just checked it again...